OpenStudy (anonymous):

A bag contains two steel balls and five brass balls. The total weight of the bag is 13 pounds. If two steel balls are added and two brass balls are removed, the bag’s weight decreases to 12 pounds. If s is the weight of a steel ball and b is the weight of a brass ball, the system of linear equations representing this situation is . _____ Solving the system of linear equations, the weight of a steel ball is ____ , and the weight of a brass ball is ____ .

OpenStudy (anonymous):

The equation is 2s+5b=13 and 4s+3b=12, For some reason when I find s & b they aren't the right numbers for their weight. I come up with fractions.. Weird ones.

OpenStudy (anonymous):

@dan815

OpenStudy (ddcamp):

Try solving to get decimal answers.

OpenStudy (anonymous):

Lol, i just noticed i seriously suck at math xD

OpenStudy (anonymous):

No one sucks, Just got to learn

OpenStudy (anonymous):

I just saw the thing and I have a headache now no joke.

OpenStudy (ddcamp):

You have the right equations, but one of the answers should be a whole number, and the other should have a decimal.

OpenStudy (dan815):

2s+5b=13 and 4s+3b=12, these are right

OpenStudy (anonymous):

s=3b/4+3 2(3b/4 +3)+5b=13 b=14/13

OpenStudy (dan815):

So lets multiply the top by -2 and add them both -2( 2s+5b=13) -4s -10b=-26 now add to this bottom equation 4s+3b=12, ----------------- -7b=-14 b=2

OpenStudy (anonymous):

Where did you get the -2

OpenStudy (dan815):

i multiplied the whole set by -2

OpenStudy (dan815):

we are allowed to multiple both sides of an equality by the same number it wont change anything

OpenStudy (dan815):

if x=2 then 2x=4 is still true

OpenStudy (dan815):

2s+5b=13 if this is true then -2( 2s+5b)=-2(13)

OpenStudy (anonymous):

.75

OpenStudy (dan815):

this is the method, they were showing you in the previous question

OpenStudy (anonymous):

1.5 I mean...

OpenStudy (anonymous):

Thanks (:

OpenStudy (dan815):

yes s=1.5 and b=2

OpenStudy (dan815):

=]

OpenStudy (anonymous):

Dont we all wish we could be geniuses